a car moves towards north at a speed of 54km/h for 1h.then iy moves eastward with same speed for same duration.the average speed and velocity of car for complete journey is ? how to find that?

Grade:12

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi ,

The total distance covered = 54 + 54 = 108 Km
The final displacement = 54^{\sqrt{}}2 Km

Average speed = Distance / time = 108 / 2 = 54 Km/h
Average velocity = Dispalcement / time = 54^{\sqrt{}}2 /2 = 27^{\sqrt{}}2 Km/h

thanks